# Inertial Focusing of Spherical Particles: The Effects of Rotational Motion

### v1

@@ Abstract
The identification of cells and particles based on their transport properties in microfluidic devices is crucial for numerous applications in biology and medicine. Neutrally buoyant particles transported in microfluidic channels, migrate laterally towards stable locations. However, the effect of the particle and flow properties on these focusing positions remain largely unknown. We conduct large scale simulations with dissipative particle dynamics, demonstrating that freely moving particles exhibit significant differences in their focusing patterns from particles that are prevented from rotation. In circular pipes, we observe drastic changes in rotating versus non-rotating focusing positions. We demonstrate that rotation-induced lateral lift force is significant, unlike previously believed, and is linearly dependent on the rotation magnitude. A simple phenomenological explanation extending existing theories is presented, that agrees well with our numerical findings. For flows in square ducts, we report four face-centered stable positions for the rotating particles, in accordance with experimental studies. However, the non-rotating particles stay scattered on a concentric one-dimensional annulus, thus revealing qualitatively different behavior with respect to the free ones. Our findings suggest new designs for micro-particle and cell sorting in microfluidics devices.
@@ 1 Introduction
In contrast to the traditional cell sorting techniques [6], such as filtering [7] or exploitation of lateral displacement [8], inertial manipulation is free from externally applied fields or direct mechanical interaction with obstacles. The first studies on inertial migration were reported in the work of Segré and Silberberg [9]. These studies reported that small particles submerged in the circular pipe flow concentrate at the annulus of radius ${\sim}0.6R$. More recently, other studies [10, 11, 12, 13] have shown that these migrations occur in channels of any cross-section shape and at a wide range of Reynolds numbers (from 10 to 1000), for rigid particles, liquid droplets and cells.
@@
Interestingly, the lateral positions where the particles focus vary significantly with Reynolds numbers, channel geometry and particle size. Matas et al. [14, 15] showed experimentally that the equilibrium particle position in circular pipe shifts outward as the channel Reynolds number increases. Choi et al. [16] confirmed that bigger particles stabilize closer to the center of the pipe. The same behavior was reported for square channels [11], with, however, an important difference: only four stable lateral positions are observed at intermediate Reynolds numbers [17], in contrast to the stable annulus in circular pipes. The migration to those points happens in two stages: during the first, faster, stage the particles gather into a deformed ring, similar to the original Segré-Silberberg annulus. During the second, slower, stage the particles move towards the four points next to the centers of the channel faces. This focusing pattern makes it possible to use the square channels to precisely collect microscale particles and cells at the channel outlet and allows for fast and non-invasive methods for sorting and filtering [11].
@@
As the small inertial effects are difficult to study analytically, various restrictions and assumptions on the problem setup were made in the theoretical works. Most notably, the particle size was assumed to be much smaller than the channel diameter [18, 19]. However, experimental verification of these findings is complicated, since small particles show very slow lateral migration, and experience pronounced Brownian motion. Recent works [20, 21] present asymptotic theory for bigger-sized particles in square channels. Their analysis captures experimental trends, however [22] displays notable inaccuracies in the near-wall region and is still limited to relatively small particles: ${\sim}0.15$ of the channel size. Overall, there is a consensus among the theoretical works that the influence of particle rotation on the lift force, and thus on the focusing position, is negligible. Numerical studies alleviate the restriction on the particle size, shape and the channel cross-section [23]. Feng [24] used two-dimensional finite element simulations to assess the forces acting on a rigid particle in Couette and Poiseuille flows, confirming that there exists a certain stable lateral position. They identified three main components of the lift force: wall lubrication, force due to the non-uniform shear rate, and force associated with the particle rotation (Magnus effect), noting once again that the rotation-attributed force is much smaller that the rest. Chun and Ladd [25] performed a three-dimensional study of the particle focusing positions in the square duct. Although they observed the two-stage migration process, their final stable points were both corners and face centers even for low Reynolds around 100, which contradicts experiments. A good agreement to the experimental data [17] was achieved by Nakagawa [22] with the immersed boundary method. They demonstrated that up to the critical channel Reynolds number of about $260$, only the face centers are stable, and after that corners become stable too. Liu [26] carried out experimental and numerical study of the particle behavior in the channel with rectangular cross-section and observed similar focusing patterns as for the square channel. They also presented results for different particle sizes, confirming that bigger particles tend to stay closer to the channel center. Harding and Stokes [27]studied the inertial focusing of neutrally buoyant spherical particles in curved microfluidic ducts at moderate Dean numbers, using regular perturbation expansions. They found that variations in the Dean number cause a change in the axial velocity profile of the background flow which influences the inertial lift force on a particle. More recently the authors studied also the effects of trapezoidal ducts [28]. In [29] simulations of particulate flows in square ducts with oblate and prolate particles revealed that inertial migration causes particles to focus in specific cross-sectional regions, with non-spherical particles occupying different positions than spherical ones. Investigations in a wavy channel [30] demonstrated that interactions between the zeroth-order lift force and the particle-free flow largely determine the focusing locations.
@@
In the present work we employ the method of dissipative particle dynamics (DPD) [31, 32, 33, 34]to study the inertial focusing of rigid particles in the circular and square channel flows. Our results are in a good agreement with experimental data and previous numerical studies. In contrast to the recent study by Huang et al. [35], we observe a broad range of Re and particle sizes where DPD is applicable (see Supplemental Material).
@@
We examine the role of particle rotation on its migration pattern in the circular pipe. We consider different particle sizes and channel Reynolds numbers. Unlike the commonly accepted assumption that rotation-induced force is negligible, we report significantly different migration of free and non-rotating particles, a behavior that is also observed in experiments [36]. According to our findings, the rotation-induced force is acting always towards the wall, in the same direction as the force produced by the non-uniform shear rate. Starting with the latter force explanation by Feng [24], we take into account flow distortion by the particle to arrive at a similar explanation for the order of magnitude and direction of the rotation-induced force. We also provide a scaling law which describes the force due to rotation in terms of the channel Reynolds number, particle size, and angular velocity. Linear dependence of this force on the rotation agrees with our explanation.
@@
In the square duct we demonstrate the mid-face equilibrium positions of the particles are reached in the previously reported two-stage process. We find that the particles with inhibited rotation show no secondary migration stage, staying scattered at the concentric annulus. We also show that unlike hypothesised by Zhou [37], the secondary migration is not solely governed by the particle rotation, but rather by more complex interplay of the rotation and wall lubrication. Our findings unveil possibilities for novel designs of microfluidic filtering devices, allowing separation of objects based on the size and rotation.
2 Simulation setup
@@
We use Mirheo, a high-throughput software for DPD microfluidic flow simulations in complex geometries with suspended bodies and viscoelastic membranes [38]. More details on the numerical method are available in the Supplemental Material.
@@
We denote the square channel side length as $H$ and for uniformity the circular pipe radius as $R=H/2$. We also introduce the particle-to-channel size ratio as $\kappa=2r/H=r/R$. The flow is driven along the channel axis by the pressure gradient $\nabla p_{ext}$. We define the channel Reynolds number as $\mathrm{Re}{}=U_{avg}\rho H/\eta$, where $U_{avg}$ is the average velocity of the undisturbed flow:
@@
To compute the forces acting on the particle at a given lateral position, we restrict the particle motion along $Oy$ and $Oz$ axes, while allowing it to freely translate with the flow along $Ox$. The instantaneous forces acting on the particle along $Oy$ and $Oz$ are random variables due to the stochastic nature of the DPD method, thus they are recorded at every time-step and then averaged. We use periodic boundary condition in $x$ and keep the length of the channel at least $15$ times larger than the particle diameter [22].
@@ 3.1 Circular pipe
In the absence of particle rotation, the equilibrium position of the particle is closer to the channel axis than in the rotating case (see Figure 2). This result agrees with previous numerical findings [24, 40]. However, we find a much more pronounced change compared to the rotating case than previously reported. We conclude from this result that the force due to rotation of the particle is larger than reported before [11, 13]. As also depicted in the Supplemental Material, the lift forces for the rotation-free particles are negatively shifted for almost all the cases and all the different particle positions $y/R$. That results in the fact that, for example, the small particle ($\kappa=0.15$) at $\mathrm{Re}{}=200$ changes its equilibrium position from the default $0.7R$ to $0.28R$.
@@
Figure 3: The lift due to the non-uniform flow shear. a: common explanations only takes into account the undisturbed flow. However, disturbed flow over the non-rotating particle (b, in blue) or freely rotating one (c, blue and red) is different, being faster close to the surface for the case of the free rotation. This leads to extra lift force attributed to the spinning (in red). d: numerical experiment shows a similar flow velocity profile relative the the particle. Simulation parameters: $\mathrm{Re}{}=100$, $\kappa=0.22$, $y/R=0.4$. The curves represent averaged stream-wise flow velocity along a line through the sphere center parallel to $Oy$, blue is non-rotating particle case, red is freely rotating one, black is undisturbed flow profile far away from the particle.
To explain the rotation significance, we have to recall the current understanding of the migration mechanism. According to numerous works [24, 6, 2], the lateral force acting on a particle can be separated into two components: the wall lubrication $F_{l}^{wall}$, which always pushes the particle away from the wall, and the force due to the non-uniform flow shear rate $F_{l}^{shear}$. While the lubrication repulsion is observed and studied in various problems with sphere translating parallel to the wall [41, 42, 43], the shear-induced lift is not understood well. Its classical explanation attributed to Feng [24] is sketched on the Figure 3a. Looking at the undisturbed flow profile relative to the moving particles, one can notice the higher and lower speed regions (shown respectively on the bottom and top of the sketch). Due to the Bernoulli-like effect, higher velocity region yields lower pressure and vice versa, thus increasing the pressure above the particle and decreasing it below. This results in a net force directed towards the wall.
@@
where linear least square fit for $y<R/2$ leads to $K\approx 5.74$. Moreover, it suggests that the rotation-induced part of the lift has the same order of magnitude as the pure $F_{l}^{shear}(0)$, since in case of the freely-rotating particles its angular velocity is reasonably close to the reference $\omega_{ref}$. More detailed force profiles available in Supplemental Material support that idea.
@@ 3.2 Square duct
For the case of the square cross-section we only vary $\mathrm{Re}{}\in\{50,100,200\}$, while keeping $\kappa=0.22$ to reduce the computational time. For the sake of brevity we hereafter focus on $\mathrm{Re}{}=100$, noting that the general findings for the other Reynolds numbers are similar according to our results. Figure 6a shows the map of lift forces acting on a freely rotating particle at different lateral positions. Due to the symmetry of the square channel, we only perform simulations for one half of the cross-section quadrant ($y>0$, $z>0$, $z\leqslant y$). For illustrative purposes, we plot forces for the full quadrant by mirroring them along the diagonal $y=z$.
@@
Figure 7: Cosine of the angle between the rotation-induced force and the angular velocity.
In the second step, we check our supposition by carrying out simulations with modified rotation of the particles on the separatrix. In these simulations, we allow the particles to freely rotate tangentially to the separatrix line, while the normal component of the rotation is modified and set to zero. In agreement with our previous conclusion, the resulting force only differs tangentially to the separatrix, while the normal component (that would have changed the location of the separatrix) stays zero. However, contrary to our initial hypothesis, the force of the secondary migration is increased, see Figure 8. Indeed, such direction of the rotation-induced force agrees with previous works [36, 24], and with our results for the circular pipe. This result means that the secondary migration is not solely attributed to the rotation, but is instead governed by the interplay of all the forces including wall lubrication, shear- and rotation-induced components. The exact physical picture could be the focus of later studies.
@@
Figure 9: When extra torque is applied to a particle the equilibrium positions change, with only a single stable edge equilibrium position apparent. Arrows are map of the lateral forces acting on a sphere in a square channel at $\mathrm{Re}{}=100$, $\kappa=0.22$ with applied external torque along $Oy$: $T^{ext}_{y}=0.75\rho U_{avg}^{2}(2r)^{5}H^{-2}$. Each arrow corresponds to one simulation and shows the direction of the force, while color marks the magnitude (red is higher). The gray line is the equilibrium manifold.

### v2

@@ Abstract
The identification of cells and particles based on their transport properties in microfluidic devices is crucial for numerous applications in biology and medicine. Neutrally buoyant particles transported in microfluidic channels, migrate laterally towards stable locations due to inertial effects. However, the effect of the particle and flow properties on these focusing positions remain largely unknown. We conduct large scale simulations with dissipative particle dynamics, demonstrating that freely moving particles exhibit significant differences in their focusing patterns from particles that are prevented from rotation. In circular pipes, we observe drastic changes in rotating versus non-rotating focusing positions. We demonstrate that rotation-induced lateral lift force is significant, unlike previously believed, and is linearly dependent on the rotation magnitude. A simple phenomenological explanation extending existing theories is presented, that agrees well with our numerical findings. In square ducts, we report four face-centered stable positions for rotating particles, in accordance with experimental studies on a range of Reynolds numbers $50\leq\mathrm{Re}\leq 200$. However, non-rotating particles stay scattered on a concentric one-dimensional annulus, revealing qualitatively different behavior with respect to the free ones. Our findings suggest new designs for micro-particle and cell sorting in inertia-based microfluidics devices.
@@ 1 Introduction
In contrast to traditional cell sorting techniques [6], such as filtering [7] or exploitation of lateral displacement [8], inertial manipulation is free from externally applied fields or direct mechanical interaction with obstacles. The first studies on inertial migration were reported in the work of Segré and Silberberg [9]. These studies reported that small particles submerged in the circular pipe flow concentrate at the annulus of radius ${\sim}0.6R$. More recently, other studies [10, 11, 12, 13] have shown that these migrations occur in channels of any cross-section shape and at a wide range of Reynolds numbers (from 10 to 1000), for rigid particles, liquid droplets and cells.
@@
Interestingly, the lateral positions where the particles focus vary significantly with Reynolds numbers, channel geometry and particle size. Matas et al. [14, 15] showed experimentally that the equilibrium particle position in circular pipe shifts outward as the channel Reynolds number increases. Choi et al. [16] confirmed that bigger particles stabilize closer to the center of the pipe. The same behavior was reported for square channels [11], with, however, an important difference: only four stable lateral positions are observed at intermediate Reynolds numbers [17], in contrast to the stable annulus in circular pipes. Migration to those points happens in two stages: during the first, faster, stage the particles gather into a deformed ring, similar to the original Segré-Silberberg annulus. During the second, slower, stage the particles move towards the four points next to the centers of the channel faces. This focusing pattern makes it possible to use the square channels to precisely collect microscale particles and cells at the channel outlet and allows for fast and non-invasive methods for sorting and filtering [11].
@@
As the small inertial effects are difficult to study analytically, various restrictions and assumptions on the problem setup were made in the theoretical works. In particular, the particle size was assumed to be much smaller than the diameter of the channel [18, 19]. However, the experimental verification of these findings is complicated, as small particles show very slow lateral migration. Recent works [20, 21] present an asymptotic theory for larger particles in square channels. Their analysis captures experimental trends; however, [22] displays notable inaccuracies in the near-wall region and is still limited to relatively small particles: ${\sim}0.15$ of the channel size. Overall, there is a consensus among the theoretical works that the influence of particle rotation on the lift force, and thus on the focusing position, is negligible. Numerical studies alleviate the restriction on the particle size, shape and the channel cross-section [23]. Feng [24] used two-dimensional finite element simulations to assess the forces acting on a rigid particle in Couette and Poiseuille flows, confirming that there exists a certain stable lateral position. They identified three main components of the lift force: wall force, force due to the non-uniform shear rate, and force associated with the particle rotation (Magnus effect), noting once again that the rotation-attributed force is much smaller than the rest. Chun and Ladd [25] performed a three-dimensional study of the particle focusing positions in the square duct. Although they observed the two-stage migration process, their final stable points were both corners and face centers even for low Reynolds around 100, which contradicts experiments. A good agreement to the experimental data [17] was achieved by Nakagawa [22] with the immersed boundary method. They demonstrated that up to the critical channel Reynolds number of about $260$, only the face centers are stable, and after that corners become stable too. Liu [26] carried out experimental and numerical study of the particle behavior in the channel with rectangular cross-section and observed similar focusing patterns as for the square channel. They also presented results for different particle sizes, confirming that bigger particles tend to stay closer to the channel center. Harding and Stokes [27] studied the inertial focusing of neutrally buoyant spherical particles in curved microfluidic ducts at moderate Dean numbers, using regular perturbation expansions. They found that variations in the Dean number cause a change in the axial velocity profile of the background flow which influences the inertial lift force on a particle. More recently, the authors also studied the effects of trapezoidal ducts [28]. In [29] simulations of particulate flows in square ducts with oblate and prolate particles revealed that inertial migration causes particles to focus in specific cross-sectional regions, with non-spherical particles occupying different positions than spherical ones. Investigations in a wavy channel [30] demonstrated that interactions between the zeroth-order lift force and the particle-free flow largely determine the focusing locations.
@@
In the present work we employ the method of dissipative particle dynamics (DPD[31, 32, 33, 34] to study the inertial focusing of rigid particles in the circular and square channel flows. Our results agree well with experimental data and previous numerical studies. In contrast to the recent study by Huang et al. [35], we observe a broad range of Re and particle sizes where DPD is applicable (see Supplemental Material [36], which also contains references [37, 38, 39, 40, 41, 42]).
@@
We investigate the influence of particle rotation on migration patterns within a circular pipe. Our study examines a range of particle sizes and channel Reynolds numbers. Contrary to the commonly held assumption that rotation-induced forces are negligible, we observe significant differences in the migration behavior of freely rotating and non-rotating particles, consistent with experimental findings [43]. Our results reveal that the rotation-induced force consistently acts toward the pipe wall, aligning with the force generated by the non-uniform shear rate. Starting with the latter force explanation by Feng [24], we take into account flow distortion by the particle to arrive at a similar explanation for the order of magnitude and direction of the rotation-induced force. We also provide a scaling law which describes the force due to rotation in terms of the channel Reynolds number, particle size, and angular velocity. Linear dependence of this force on the rotation agrees with our explanation.
@@
In the square duct we demonstrate the mid-face equilibrium positions of the particles are reached in the previously reported two-stage process. We find that the particles with inhibited rotation show no secondary migration stage, staying scattered at the concentric annulus. We also show that unlike hypothesised by Zhou [44], the secondary migration is not solely governed by the particle rotation, but rather by more complex interplay of the rotation and wall force. Our findings unveil possibilities for novel designs of microfluidic filtering devices, allowing separation of objects based on the size and rotation.
2 Methods
We consider straight pipes of circular and square cross sections filled with a Newtonian, incompressible fluid, with periodic boundary conditions to mimic long pipes. We model the suspended particle as a spherical, non-deformable object with the same density as the fluid. To simulate this system, we employ the DPD method within the software Mirheo [45], which has been extensively validated for incompressible flows with suspended biological cells [46, 47], and suspended rigid objects [45, 48]. In addition, several cases of particles in circular pipes and square ducts are validated against experimental work [16, 17] and numerical work [11, 26, 22] in the Supplemental Material [36]. The parameters of the numerical model that were used in this work are reported in the Supplemental Material [36].
@@ 2 Methods
We denote the square channel side length as $H$ and for uniformity the circular pipe radius as $R=H/2$. We also introduce the particle-to-channel size ratio as $\kappa=2r/H=r/R$. The flow is driven along the channel axis by the pressure gradient $\nabla p_{ext}$. We define the channel Reynolds number as $\mathrm{Re}{}=U_{avg}\rho H/\eta$, where $U_{avg}$ is the average velocity of the undisturbed flow,
@@
Furthermore we define the particle Reynolds number as $\mathrm{Re}_{p}=\kappa^{2}\mathrm{Re}$ [24]. To compute the forces acting on the particle at a given lateral position, we restrict the particle motion along $Oy$ and $Oz$ axes, while allowing it to freely translate with the flow along $Ox$. The instantaneous forces acting on the particle along $Oy$ and $Oz$ are noisy due to the stochastic nature of the DPD method. Therefore, we average these forces over time after the simulation has equilibrated. The simulation is considered at equilibrium after the velocity of the particle along $x$ reached a constant value. We use periodic boundary condition in $x$ and keep the length of the channel at least $15$ times larger than the particle diameter [22].
@@ 3.1 Circular pipe
In the absence of particle rotation, the equilibrium position of the particle is closer to the channel axis than in the rotating case (see Figure 2). This result agrees with previous numerical findings [24, 50]. However, we find a much more pronounced change compared to the rotating case than previously reported. We conclude from this result that the force due to rotation of the particle is larger than reported before [11, 13]. As also depicted in the Supplemental Material [36], the lift forces for the rotation-free particles are negatively shifted for almost all the cases and all the different particle positions $y/R$. That results in the fact that, for example, the small particle ($\kappa=0.15$) at $\mathrm{Re}{}=200$ ($\mathrm{Re}_{p}=4.5$) changes its equilibrium position from the default $0.7R$ to $0.28R$.
@@
Figure 4: The lift due to the non-uniform flow shear. a: common explanations only takes into account the undisturbed flow. However, disturbed flow over the non-rotating particle (b, in blue) or freely rotating one (c, blue and red) is different, being faster close to the surface for the case of the free rotation. This leads to extra lift force attributed to the spinning (in red). d: numerical experiment shows a similar flow velocity profile relative the the particle. Simulation parameters: $\mathrm{Re}{}=100$, $\kappa=0.22$, $y/R=0.4$ ($\mathrm{Re}_{p}=4.84$). The curves represent averaged stream-wise flow velocity along a line through the sphere center parallel to $Oy$, blue is non-rotating particle case, red is freely rotating one, black is undisturbed flow profile far away from the particle.
To explain the rotation significance, we have to recall the current understanding of the migration mechanism. According to numerous works [24, 6, 2], the lateral force acting on a particle can be separated into two components: the wall force $F_{l}^{wall}$, which always pushes the particle away from the wall, and the force due to the non-uniform flow shear rate $F_{l}^{shear}$. While the wall repulsion is observed and studied in various problems with sphere translating parallel to the wall [51, 52, 53], the shear-induced lift is not understood well. Its classical explanation attributed to Feng [24] is sketched on the Figure 4a. Looking at the undisturbed flow profile relative to the moving particles, one can notice the higher and lower speed regions (shown respectively on the bottom and top of the sketch). We consider two streamlines that are on each side of the particle but close to each other upstream, where they share the same velocity and pressure. Thus, applying the Bernoulli principle along each streamline, the difference in velocity on each side of the particle yields a larger pressure above the particle than below it [24]. This results in a net force directed towards the wall.
@@
where linear least square fit for $y<R/2$ leads to $K\approx 5.74$. Moreover, it suggests that the rotation-induced part of the lift has the same order of magnitude as the pure $F_{l}^{shear}(0)$, since in case of the freely-rotating particles its angular velocity is reasonably close to the reference $\omega_{ref}$. More detailed force profiles available in Supplemental Material [36] support that idea.
@@ 3.2 Square duct
For the case of the square cross-section we only vary $\mathrm{Re}{}\in\{50,100,200\}$, while keeping $\kappa=0.22$ to reduce the co mputational time. This corresponds to particle Reynolds numbers $\mathrm{Re}_{p}\in\{2.42,4.84,9.68\}$. For the sake of brevity we hereafter focus on $\mathrm{Re}{}=100$, noting that the general findings for the other Reynolds numbers are similar according to our results. Figure 6a shows the map of lift forces acting on a freely rotating particle at different lateral positions. Due to the symmetry of the square channel, we only perform simulations for one half of the cross-section quadrant ($y>0$, $z>0$, $z\leqslant y$). For illustrative purposes, we plot forces for the full quadrant by mirroring them along the diagonal $y=z$.
@@
Figure 7: Heat map corresponds to the dot product of unit vectors in the direction of the rotation-induced force $F_{l}^{rot}$ and angular velocity $\omega$, representing their alignment.
In the second step, we check our supposition by carrying out simulations with modified rotation of the particles on the separatrix. In these simulations, we allow the particles to freely rotate tangentially to the separatrix line, while the normal component of the rotation is modified and set to zero. In agreement with our previous conclusion, the resulting force only differs tangentially to the separatrix, while the normal component (that would have changed the location of the separatrix) stays zero. However, contrary to our initial hypothesis, the force of the secondary migration is increased, see Figure 8. Indeed, such direction of the rotation-induced force agrees with previous works [43, 24], and with our results for the circular pipe. This result means that the secondary migration is not solely attributed to the rotation, but is instead governed by the interplay of all the forces including wall force, shear- and rotation-induced components. The exact physical picture could be the focus of later studies.
@@
Figure 9: When extra torque is applied to a particle the equilibrium positions change, with only a single stable edge equilibrium position apparent. Arrows are map of the lateral forces acting on a sphere in a square channel at $\mathrm{Re}{}=100$, $\kappa=0.22$ ($\mathrm{Re}_{p}=4.84$) with applied external torque along $Oy$: $T^{ext}_{y}=0.75\rho U_{avg}^{2}(2r)^{5}H^{-2}$. Each arrow corresponds to one simulation and shows the direction of the force, while color marks the magnitude (red is higher). The gray line is the equilibrium manifold.